Two factory plants are making TV panels. Yesterday, Plant A produced 8,000 panels. 2 % of the panels from Plant A and 6% of the panels from Plant B were defective. How many panels did Plant B produce, if the overall % of defective panels from the two plants was 5%?
Answer by jim_thompson5910(35256)

"Plant A produced 8,000 panels. 2 % of the panels from Plant A...were defective"

8000*0.02 = 160 panels from plant A are defective.

-----------------------------------------------------------------------

Let x = number of panels plant B produces

"6% of the panels from Plant B were defective"

There are 0.06x defective panels that come from plant B

------------------------------------------------------------------------

In total, we have 160+0.06x panels that are defective.

This is out of 8000+x panels total.

Their ratio must equal 0.05 since " the overall % of defective panels from the two plants was 5%"

So...



(160+0.06x)/(8000+x) = 0.05

160+0.06x = 0.05(8000+x)

160+0.06x = 0.05(8000)+0.05(x)

160+0.06x = 400+0.05x

160+0.06x-0.05x = 400

0.06x-0.05x = 400-160

0.01x = 240

x = 240/0.01

x = 24000




So that means plant B has produced 24,000 panels.

Of these panels, 6% are defective. So 0.06*24000 = 1440 are defective.

-------------------------------------------------------

Notice how 160+1440 = 1600 total defective panels out of 8000+24000 = 32000 gives you

1600/32000 = 0.05

which reflects the overall defect percentage of 5%